OpenStudy (anonymous):

In the following equation, identify the x-intercepts in the graph y = 3x2 + 19x – 40The dimensions of Eric’s pool are 9 feet by 7 feet. He wants to build a walkway surrounding the pool so that the square footage of the walkway and pool is 143 square feet. How wide is one side of the walkway?

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

Let x = the width of the pool Since the walkway goes all the way around the pool the over dimensions will be: (2x+9) by (2x+7); then the overall area (pool & walkway) will be: (2x+9)*(2x+7) = 143 FOIL 4x^2 + 14x + 18x + 63 = 143 4x^2 + 32x + 63 - 143 = 0 4x^2 + 32x - 80 = 0; quadratic equation Simplify, divide by 4 x^2 + 8x - 20 = 0 Factors to (x+10)(x-2) = 0 The positive solution is what we want here x = 2 ft is the width of the walkway